Circular No. 6217

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1995 Q2

R. H. McNaught, Anglo-Australian Observatory, reports the discovery of a comet by Malcolm Hartley on a film taken by Michael J. Drinkwater. The following precise positions are available:

1995 UT α_{2000} δ_{2000} m_{1} Observer
Aug. 29.54634 22^{h}19^{m}56.^{s}11 −24^{o}30′10′′.1 14 Drinkwater
29.60190 22 19 26.48 −24 31 54.6  _{”}
30.31385 22 13 05.10 −24 54 03.3  Balam
30.31543 22 13 04.27 −24 54 05.9  _{”}
30.31697 22 13 03.53 −24 54 08.9  _{”}
30.36163 22 12 39.56 −24 55 29.1  _{”}
30.36315 22 12 38.82 −24 55 32.0  _{”}
30.36471 22 12 38.04 −24 55 35.5  _{”}
30.39038 22 12 24.88 −24 56 07.9  Ryan
30.39127 22 12 24.41 −24 56 09.2  _{”}
30.39216 22 12 23.97 −24 56 11.0  _{”}
30.39305 22 12 23.53 −24 56 13.1  _{”}
30.39394 22 12 22.99 −24 56 14.5  _{”}
30.42096 22 12 08.53 −24 57 03.2  Garradd
30.42661 22 12 05.51 −24 57 11.8  _{”}
30.43525 22 12 00.86 −24 57 29.2  _{”}

M. J. Drinkwater (Siding Spring). 1.2-m U.K. Schmidt. Measurer R. H. McNaught. Comet very condensed, 5′′ coma, no tail. Faint image, beginning of trail involved with star.

D. D. Balam (Climenhaga Observatory). 0.5-m reflector + CCD. S. G. Ryan (Siding Spring). Australian National Observatory 1.0-m reflec-

tor + CCD. Measurer R. H. McNaught. Analysis by M. J. Drinkwater shows a core of FWHM 3′′.9 measured in seeing of FWHM 2′′.2, a coma of full diameter 10′′–15′′ at the sky limit and a tail extending 20′′ or more at p.a. 45^{o}.

G. Garradd (Loomberah). 0.25-m f/4.1 reflector + CCD. Weak images through patchy cloud. m_{1} = 13 visually at 80× on Aug. 30.44.

COMET 81P/WILD 2

A. Fitzsimmons and M. Cartright, Queen’s University, Belfast, observing with the 1-m Kapteyn reflector at La Palma, report this comet as of near-stellar appearance at R = 22 on Aug. 26.20 and 27.13 UT.

1995 August 30 Brian G. Marsden